a) A wind instrument is designed to be a closed tube that is 80 cm long. With a speed of sound of 340 m/s, what is the frequency required to get the instrument to resonate in its 3rd harmonic? Round your answer to 3 significant digits.
b) The tube is then opened at the other end. Assuming the speed of sound has not changed, what would the frequency need to be to resonate in its 3rd harmonic now? Round your answer to 3 significant digits
1
Expert's answer
2020-06-04T09:01:01-0400

a)


"f=\frac{3v}{4L}=\frac{3(340)}{4(0.8)}=319\ Hz"

b)


"f=\frac{3v}{2L}=\frac{3(340)}{2(0.8)}=638\ Hz"